OpenStudy (adrianna.gongora):

Shannon is managing a small zoo and has been analyzing the attendance data. Shannon has found that the number of visitors increases exponentially as the temperature increases. Shannon has also found a linear equation that models the number of people who leave the park early depending on the temperature. Describe how Shannon can combine these two functions into a new function and explain what that function would predict.

OpenStudy (adrianna.gongora):

@wolf1728

OpenStudy (wolf1728):

That seems like a complex question. First, if attendance increases exponentially as temperature increases you could have a formula such as attendance (people) = 10^((Temp -70)/10) So at 70 degrees, you'd have 1 person 10^((70-70)/10) = 10^0 = 1 At 80 degrees, you'd have 10 people 10^((80-70)/10) = 10^(10/10) = 10^1 = 10 90 degrees = 100 people That seems like a formula that needs some fine tuning because 100 degrees = 1,000 people 110 degres = 10,000 people !!!

OpenStudy (adrianna.gongora):

so if the degrees go up 10 then theres a 0 added to the people?

OpenStudy (wolf1728):

Yes, but that is just a formula I invented, If nothing else, the question should set more parameters. I think you might have to get someone else to help here or see if the instructor / or the book has some further suggestions.

OpenStudy (adrianna.gongora):

okay thank you.

OpenStudy (wolf1728):

u r welcome :-)